What is the molarity of the resulting solution when 47.946 g of Mn(ClO4)2 · 6 H2O are added to 250.0 mL of water?
Answer #1

Molar mass of Mn(ClO4)2.6H2O = 55 + (35.5 + 64)2 + 6(16+1+1) = 362 gm/mol

Number of moles of Mn(ClO4)2. 6H2O = 47.946/362 = 0.13244 moles

Molarity = Number of moles of Solute/Volume of Solution (in L)

=> 0.13244/ (250/1000)

=> 0.5297 M